It takes a contractor 4 days to paint a house. His apprentice takes 6 days to paint the
same house. If the contractor and his apprentice paint the house together, how many
days will it take the both of them to complete painting 1 house?

Answer by richwmiller(17219)

x/4+x/6=1
6x+4x=24
10x=24
x=2.4 days
each day they can do 1/4 and 1/6
6/24+4/24=
10/24 =5/12